OpenStudy (anonymous):

please help me figure out why 7r= -7/2(thats a fraction) the answer is 1/2 but why and how do i show my work to get this answer?

OpenStudy (lgbasallote):

7r = -7/2 doesnt make 1/2 o.O are you sure you wrote the question right?

OpenStudy (anonymous):

its in my practice test which shows the answer is -1/2 but i got 3.5

OpenStudy (lgbasallote):

ahh it's -1/2 right

OpenStudy (anonymous):

they want me to show my work but i dont understand how its negative half

OpenStudy (lgbasallote):

\[7r = -\frac{7}{2}\] divide both sides by seven \[\frac{\cancel{7}r}{\cancel{7}} = -\frac{\cancel{7}}{2} \times \frac{1}{\cancel{7}}\] \[r = -\frac{1}{2}\] got it now?

OpenStudy (anonymous):

where did you get the 1/7 from?

OpenStudy (lgbasallote):

divide by 7

OpenStudy (lgbasallote):

that also means multiply by 1/7

OpenStudy (lgbasallote):

\[\frac{a}{7} = a \times \frac{1}{7}\]

OpenStudy (anonymous):

and quick question how are you able to type on the keyboard the fraction correctly without using the slash?

OpenStudy (lgbasallote):

type in your equation editor: \frac{numerator}{denominator} this produces \(\frac{numerator}{denominator}\)

OpenStudy (lgbasallote):

you get it right?

OpenStudy (anonymous):

yea i was multiplying straight across instead of cross multiply

OpenStudy (lgbasallote):

great!! ^_^